# Einstein’s basement - a model for dark matter and an expanding universe?

### v1

@@ IV Summary, Outlook and Conclusions
If our model can survive the comparison with astrophysical observations there would be many more options to explore new physics. In this paper we only considered classical particle dynamics and compared our results to those of classical Newtonian physics. However, with the full Hamiltonian (I) at hand, an extension of our analysis towards a comparison with special and general relativity is possible. This is of particular interest, since the concept of relativity does not hold for the sector of Einstein’s basement, while it is still preserved for regular matter in our model. However, considering interactions between particles of both sectors, violations of the relativity principle, such as a breaking of local Lorentz invariance, can be expected also for regular matter.
@@
A field theoretical approach of our model could elucidate the nature of the field whose excitations are the quasi particles, we identified as regular particles and particles from Einstein’s basement. If we stress the similarities of our model with the description of polaritons in a crystal [20, 21, 22], the underlying field would correspond to the crystal lattice in that case. In particular, the relation of the mass or energy density of this hypothetical background field to concepts like the cosmological constant and dark energy has to be investigated.
@@
A consequent extension of the work presented here is to study the given EMRs in a quantum mechanical framework and, ultimately, within the formalism of quantum field theory. Particularly interesting is the regime of small energies and momenta, where all EMRs in Einstein’s basement closely approach each other and the photon line.
@@ Acknowledgements.
We thank Uwe Sterr, Piet Schmidt, Andrey Surzhykov, Eberhard Tiemann and Jun Ye for fruitful discussions and Christopher Mertens for numerical simulations. We acknowledge the support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy—EXC 2123 QuantumFrontiers—390837967. FR thanks QUEST Leibniz Research School for longstanding support.

### v2

@@ V Summary, Conclusions and Outlook
We have complemented the EMRSR of special relativity by a second EMREB in Einstein’s basement derived from a forbidden crossing between the constant rest mass of a particle and a particle with energy $pc$. As a result, regular matter and matter originating from Einstein’s basement are considered as made of quasi particles of a yet unknown field. We studied the kinematics of the newly introduced particles and found that their velocity $v$ is replaced by $c-v$ in comparison to the regular case. In this paper, we only considered classical particle dynamics and compared our results to those of classical Newtonian physics. However, with the full Lagrangian (IV) at hand, an extension of our analysis towards a comparison with special and general relativity is possible. This is of particular interest, since the concept of relativity does not hold for the sector of Einstein’s basement, while it is still preserved for regular matter in our model. Considering interactions between particles of both sectors, violations of the relativity principle, such as a breaking of local Lorentz invariance, can be expected also for regular matter. In theory, extensions of the standard model of particles have been proposed which include general relativity and many possible operators that break Lorentz symmetry [36]. Investigations of such potential violations are ongoing both theoretically [37] and experimentally with different means in pulsars [38, 39], ions [40], atom interferometers [41] and others.
Assuming that Newtonian gravity also holds for the quasi particles in Einstein’s basement, we found that under certain conditions also repulsion between the particles can occur. Furthermore, regular particles are dragged along with basement particles. Combining these two findings, this mechanism can describe an expansion of a mixed system in space. Whether this hypothesis can contribute or explain the observed accelerated expansion of the universe, as it is suggested by the study of type Ia supernovae candles [42, 43], needs to be investigated.
Einstein’s basement might also be a region where the long searched for constituents of dark matter could hide. Properties of dark matter particles have been investigated [44, 10] and a number of theoretical hypotheses has been offered for the origin of dark particles like weakly interacting massive particles (WIMPs) [45], axion-like particles [46], dark photons [47], erebons [48], Kaluza-Klein particles [49] or others. However, none of these particles have been found experimentally up to now [50, 51]. A possible reason might be that the searched particles may obey a different kinematics, such as the one described in our model.
The existence of two branches of different energies (Fig. 1) suggests diabatic transitions under a suitable interaction between the two branches. Diabatic electromagnetic transitions are well known to occur in the case of an avoided crossing. Such transitions were modeled by Landau [52], Zener [53], Stückelberg [54] and Majorana [55] and have been tested experimentally in quantum optics [56]. Assuming that transitions under the influence of a particular interaction are allowed, they would lead to a decrease of the population in the $E_{SR}$ state and increase the population in the $E_{EB}$ state. However, since no signals of such transitions have been observed so far, we conclude a very low probability for the decays to occur and can exclude electroweak and strong interaction as initiating such a decay. Considering interactions of this kind unlikely, particles from both sectors may be still allowed to interact gravitationally by virtue of their positive energy densities.
A future task is to checked, whether our model delivers an competitive explanation for the existing measurement data and provides an alternative to the current maps of the dark matter distribution in the galaxies [57, 58]. Revisiting the huge number of existing data and the new data expected from the EUCLID space mission [59] could give a hint whether the presented ideas should be pursued further. There is furthermore to be checked whether the idea of Einstein’s basement can be applied at macroscopic, i.e., astrophysical or cosmological scales or whether it can provide meaningful interpretations only at the scales of quantum physics, such as it is the case for Dirac theory.
@@
A field theoretical approach of our model could elucidate the nature of the field whose excitations are the quasi particles, we identified as regular particles and particles from Einstein’s basement. If we stress the similarities of our model with the description of polaritons in a crystal [2, 3, 4], the underlying field would correspond to the crystal lattice in that case. In particular, the relation of the mass or energy density of this hypothetical background field to concepts like the cosmological constant and dark energy might be a promising investigation.
A consequent extension of the work presented here is to study the given EMRs in a quantum mechanical framework and, ultimately, within the formalism of quantum field theory. Since the inner symmetries of the fields, such as the $U(1)$-gauge symmetry, exist independently of the considered group of motion, it is straightforward to write a theory of gauge fields, either in special relativity or in Einstein’s basement.
@@
Particularly interesting is the regime of small energies and momenta, where all EMRs in Einstein’s basement closely approach each other near the $pc$-line.
@@ Acknowledgements.
We thank Uwe Sterr, Piet Schmidt, Andrey Surzhykov, Eberhard Tiemann and Jun Ye for fruitful discussions and Christopher Mertens for numerical simulations and Hauke Uhde for his analysis of the implications of our model for galaxy dynamics (work in progress). We acknowledge the support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy—EXC 2123 QuantumFrontiers—390837967. FR thanks QUEST Leibniz Research School for longstanding support.